OpenStudy (anonymous):

What is the degrees Fahrenheit if the temperature is -60°C? Show all work. Remember to label your answer. If necessary, round your final answer to the nearest tenth.

OpenStudy (31356):

-76F

OpenStudy (31356):

Here's How: Take the temperature in Fahrenheit subtract 32. Divide by 1.8. The result is degrees Celsius.

OpenStudy (31356):

Medal?

OpenStudy (mathmale):

Love (referring to Homeworked, not to 0031356), there's a formula relating temperature in degrees Centigrade to temperature in degrees Fahrenheit. Do you know this formula?

OpenStudy (anonymous):

Yes love, i know the answer but im not sure how to exactly show my work.

OpenStudy (mathmale):

If not, one way of obtaining that formula would be to do an Internet search for "conversion of degrees Fahrenheit to degrees Celcius."

OpenStudy (31356):

Thanks for the medal!

OpenStudy (mathmale):

You want to convert -60 degrees Celsius to degrees Fahrenheit. There is also a formula for conversion of temperature in degrees C to degrees F; Do you happen to have that in front of you?

OpenStudy (anonymous):

Love, i know the answer but im not show how to show my work on paper.

OpenStudy (anonymous):

because the answer is -76F correct?

OpenStudy (mathmale):

Conversion of F to C: C = (9/5)(F-32). Conversion of C to F: F = 5C/9 + 32. Although I won't insist that y ou have to do this, I do recommend it. So, to convert -60 de. C to deg. F, substitute c=-60 into the 2nd equation. Let me check, but I'm still gonna type out every step: F = 5(-60)/9 + 32 F = -300/9 + 32 F = -100/3 + 32 F= -100/3 + 3(32)/3 (so that both quantities have the same denominator) F = -99/3 = -30 degrees F. I'm going to ask YOU to check MY math; it's worth the effort to do so.

OpenStudy (anonymous):

Hm? The answer is -76F love..

OpenStudy (mathmale):

Hint: my work is wrong. Let me ask you to evaluate this: -100/3 Then, evaluate -100/3 + 32.

OpenStudy (mathmale):

-100/3=?

OpenStudy (mathmale):

My Sweet??

OpenStudy (anonymous):

Give me a sec!

OpenStudy (mathmale):

I've given you a whole bunch of secs already! :)

OpenStudy (anonymous):

im multitasking right now!

OpenStudy (mathmale):

Aha. Thought so. I do that quite often myself. But I've got to get off OpenStudy mighty soon.

OpenStudy (mathmale):

-100/3 = -33 1/3. What is -100/3 + 32? Love?

OpenStudy (anonymous):

Waiiit.

OpenStudy (anonymous):

I thought it would be -33.(3333333)

OpenStudy (mathmale):

I note that in this problem you need to convert FROM degrees C TO degrees F. The approach that our friend 0031356 used is valid for conversions in the opposite direction: FROM F to C. Yes. -33.3333 + 32 = ?

OpenStudy (mathmale):

-33.3333 +32 --------- ? So, -60C => -4/3 degrees F.

OpenStudy (anonymous):

right, thats what i thought.. so.. 1.3333333 then?

OpenStudy (anonymous):

-1.3333333*

OpenStudy (mathmale):

What do YOU think? you insisted it was -76. Try to decide which result makes the more sense.

OpenStudy (anonymous):

RIght.. thanks love

OpenStudy (mathmale):

Great working with you. Have to get off Open Study now. See you soon, love. MM